2 Al + 3 FeSO3 -> Al2(SO3)3 + 3 Fe


if 136 grams of aluminum are consumed, how many grams of iron are produced?


1
Expert's answer
2021-06-10T03:31:53-0400

When 2 mole of Al ( Aluminum) are consumed 3 moles of Fe ( Iron ) produced is 2:3

136 grams of aluminum means 136/26.98 = 5.04 moles

So moles of Iron produces from Aluminum is 7.54

So mass of Fe produces = 7.54× 55.84 = 422.15g